# Deciphering the Scattering of Polydisperse Hard Spheres using Deep Learning

### v1

@@ I Introduction
Despite the fundamental importance, accurately modeling the scattering function of polydisperse hard spheres, especially at higher concentrations and polydispersity, has been difficult using conventional analytical theories. Traditional approaches, such as the Percus-Yevick approximation[15, 16] for the structure factor combined with a form factor averaging scheme[4] ($\beta$ correction for polydispersity), attempt to decouple the particle shape and inter-particle correlations. While these methods offer valuable insights for dilute or monodisperse systems, their fundamental assumptions break down under conditions of increasing volume fraction and polydispersity[17, 11]. This oversight leads to significant discrepancies between theoretical solutions and experimental or simulation data, limiting their predictive power for scattering data analysis. Consequently, there is a pressing need for more robust and accurate methods that can capture the complex scattering behavior of polydisperse systems.
Machine learning (ML)[18, 19], and deep learning[20, 21], has transformed data analysis in science by uncovering complex pattern in large datasets. In soft matter physics and scattering analysis, ML models have been applied to analyze scattering function of systems like colloids[22, 23], lamellar[24, 25], and polymer[26, 27, 28, 29] systems. These ML models utilize simulation data to learn the mapping between the scattering function and the corresponding system parameters. This data-driven, simulation-based approach[30, 22] offers a powerful alternative to traditional analytical method, motivating their use for interpreting the scattering data of interacting polydisperse colloids.
@@ II.1 Molecular Dynamics
To model the polydispersity of the hard spheres, we consider three kinds of size distribution that commonly used and available in existing software packages such as SASView: uniform distribution, normal distribution, and lognormal distribution. We use $D_{0}$ to indicate the mean diameter for the uniform and normal distribution, and the median diameter of the lognormal distribution, $D_{0}=1$ is also chosen as the natural unit in the simulation.
II.2 Scattering analysis
@@
[…]
@@
[…]
@@
[…]
@@ III.1 Scattering function of hard spheres
We then prepare the a dataset consisting of $5,000$ $(\eta,\sigma)$ and corresponding scattering function $I(Q)$ by running the MD simulation for the uniform-distribution polydisperse hard spheres system. The volume fraction and polydispersity are randomly selected from $\eta\in U(0,0.5)$ and $\sigma\in U(0,0.3)$, independently. And the scattering function $I(Q)$ are calculated for structure vector $Q$ uniformly placed on the $100$ grid points within $Q\in[3,13]$. Combining all of the scattering function into $\mathbf{F}=\{\log_{10}{I(Q)}\}$, which is a $5000\times 100$ matrix. To analyze the dataset $\mathbf{F}$, we carry out principle component analysis[44] by calculating the singular value decomposition[45] of the dataset $\mathbf{F}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{T}$, such that $\mathbf{U}$ is $5000\times 5000$, $\mathbf{\Sigma}$ is $5000\times 100$, $\mathbf{V}$ is $100\times 100$, and the diagonal entries of $\mathbf{\Sigma}$ are the singular values of $\mathbf{F}$, determining the coefficient of the projection of $\mathbf{F}$ onto $\mathbf{V}$.
@@ III.4 Other distributions
Figure 9: Illustration of the three different size distribution of 256 hard spheres with volume fraction $\eta=0.1$. (a) Uniform distribution in $D\sim U[D_{0}-\sigma,D_{0}+\sigma]$ with $D_{0}=1$ and $\sigma=0.1$. (b) Normal distribution $D\sim\mathcal{N}(D_{0},\sigma)$ with mean $D_{0}=1$ and standard deviation $\sigma=0.1$. (c) Log-normal distribution $\log D\sim\mathcal{N}(\log{D_{0}},\sigma)$ ) with median $D_{0}=1$ and $\sigma=0.1$. (d) Probability density function of three kinds of size distribution.
While the above detailed study are mainly focused on one type of particle size distribution, our approach is general, and can easily apply to any other polydispersity distribution. The commonly used ones are uniform, normal, and lognormal. Fig. 9 shows sample configurations of the hard spheres and the diameter distribution of these three distribution. The diameter of the hard sphere particles are sampled from the corresponding distribution, parameterized by the polydispersity $\sigma$. For instance, the uniform distribution results in an uniform distribution of the particle diameter $D$ in the continuous interval $[D_{0}-\sigma,D_{0}+\sigma]$ and we use $D_{0}=1$ as the natural unit. For the normal distribution, the diameter $D$ of each particle follow the normal distribution with mean $D_{0}=1$ and standard deviation $\sigma$, such that $D\sim\mathcal{N}(D_{0},\sigma)$. Finally, for the lognormal distribution, the log of the particle diameter follows the normal distribution such that $\log D\sim\mathcal{N}(\log{D_{0}},\sigma)$ ), the $D_{0}=1$ become the median diameter.
@@ IV Summary
This work lays the groundwork for applying ML-based analysis to the scattering of polydisperse systems. Moving forward, this framework can be applied to even more intricate systems, such as charged colloids, binary colloids[46], rods and charged rods[47], as well as other dense dispersion systems, where the strong coupling between the form factor and structure factor renders traditional analytic modeling ineffective. Since polydispersity is ubiquitous in nature, extending our framework to these cases will greatly improve the analytical power of the scattering technique.
@@ Acknowledgements.
LD thanks Jan-Michael Carrillo for fruitful discussion. This research used resources at the Spallation Neutron Source and the Center for Nanophase Materials Sciences, US Department of Energy (DOE) Office of Science User Facilities operated by Oak Ridge National Laboratory. This research was sponsored by the Laboratory Directed Research and Development Program of Oak Ridge National Laboratory, managed by UT-Battelle, LLC, for the US DOE. Computations used resources of the Oak Ridge Leadership Computing Facility, which is supported by the DOE Office of Science under contract No. DE-AC05-00OR22725. Application of machine learning to soft matter was supported by the US DOE, Office of Science, Office of Basic Energy Sciences Data, Artificial Intelligence and Machine Learning at DOE Scientific User Facilities Program under award No. 34532.

### v2

# Deciphering the Small-Angle Scattering of Polydisperse Hard Spheres using Deep Learning

@@ I Introduction
Despite the fundamental importance, accurately modeling the scattering function of polydisperse hard spheres, especially at higher concentrations and polydispersity, has been difficult using conventional analytical theories. Traditional approaches, such as the Percus-Yevick approximation[15, 16] for the structure factor combined with a form factor averaging scheme[4] ($\beta$ correction for polydispersity), attempt to decouple the particle shape and inter-particle correlations. While these methods offer valuable insights for dilute or monodisperse systems, their fundamental assumptions break down under conditions of increasing volume fraction and polydispersity[17, 11, 18, 19]. This oversight leads to significant discrepancies between theoretical solutions and experimental or simulation data, limiting their predictive power for scattering data analysis. Consequently, there is a pressing need for more robust and accurate methods that can capture the complex scattering behavior of polydisperse systems.
Machine learning (ML)[20, 21], and deep learning[22, 23], has transformed data analysis in science by uncovering complex pattern in large datasets. In soft matter physics and scattering analysis, ML models have been applied to analyze scattering function of systems like colloids[24, 25, 26], lamellar[27, 28], and polymer[29, 30, 31, 32] systems. These ML models utilize simulation data to learn the mapping between the scattering function and the corresponding system parameters. This data-driven, simulation-based approach[33, 24] offers a powerful alternative to traditional analytical method, motivating their use for interpreting the scattering data of interacting polydisperse colloids.
@@ II.1 Molecular Dynamics
To model the polydispersity of the hard spheres, we consider three kinds of size distribution that commonly used and available in existing software packages such as SASView, which is widely adopted by the small angle scattering community: uniform distribution, normal distribution, and lognormal distribution. We use $D_{0}$ to indicate the mean diameter for the uniform and normal distribution, and the median diameter of the lognormal distribution, $D_{0}=1$ is also chosen as the natural unit in the simulation.
II.2 Small-angle scattering analysis
@@
where $\left$$\dots\right$> is the ensemble average of all configurations, $\mathbf{Q}$ is the scattering vector, $\mathbf{r}_{i}$ is the position of particle $i$, $D_{i}$ is the diameter of particle $i$, and $F(Q;D_{i})$ is the form factor amplitude of the hard sphere, which is given by[41]:
@@
in which $S_{PY}(Q)$ is the Percus-Yevick approximation of the hard spheres structure factor which depend on the volume fraction $\eta$, and $P(Q)$ is the form factor given by $P(Q)=\left<F(Q;D)^{2}\right>_{D}/\left<(\frac{\pi D^{3}}{6})^{2}\right>_{D}$. To account for polydispersity of hard spheres, a $\beta$ correction [4] was introduced, such that
@@
where $\beta=\left<F(Q;D)\right>_{D}^{2}/\left<F(Q;D)^{2}\right>_{D}$. The detailed calculation of $S_{PY}(Q)$ can be found in the corresponding references[15, 16, 43]. Apparently, the traditional method oversees the coupling between the form factor and inter-particle structure factor, and we expect such approximation to fail at higher volume fraction where the interaction between the particles start to reflect on the scattering function. It is crucial to note that the scattering analysis methods described here are not applicable to optical or light scattering methods[44], where different physical principles and near-field effects need to be considered.
@@ III.1 Scattering function of hard spheres
We then prepare the a dataset consisting of $5,000$ $(\eta,\sigma)$ and corresponding scattering function $I(Q)$ by running the MD simulation for the uniform-distribution polydisperse hard spheres system. The volume fraction and polydispersity are randomly selected from $\eta\in U(0,0.5)$ and $\sigma\in U(0,0.3)$, independently. And the scattering function $I(Q)$ are calculated for structure vector $Q$ uniformly placed on the $100$ grid points within $Q\in[3,13]$. The $Q$ range is chosen such that we focus on the interaction peak at low $Q$ and also cover the first first dip of sphere form factor. High order minima are not taken into consideration since there is no interaction effect at these $Q$ range. Combining all of the scattering function into $\mathbf{F}=\{\log_{10}{I(Q)}\}$, which is a $5000\times 100$ matrix. To analyze the dataset $\mathbf{F}$, we carry out principle component analysis[48] by calculating the singular value decomposition[49] of the dataset $\mathbf{F}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{T}$, such that $\mathbf{U}$ is $5000\times 5000$, $\mathbf{\Sigma}$ is $5000\times 100$, $\mathbf{V}$ is $100\times 100$, and the diagonal entries of $\mathbf{\Sigma}$ are the singular values of $\mathbf{F}$, determining the coefficient of the projection of $\mathbf{F}$ onto $\mathbf{V}$.
@@ III.4 Other distributions
[…]
@@ IV Summary
This work lays the groundwork for applying ML-based analysis to the scattering of polydisperse systems. Moving forward, our approach can be applied on experimental data, and both the generator and inferrer can be used under different conditions. When using the inferrer, one needs to firstly find the mean particle size for data normalization using dilute sample, then can just apply the inferrer for sample of higher concentration to achieve one-pass inference. When there is not dilute sample for reference or the instrument resolution is not high enough, one can combing the generator with a regression model to fit the experimental data[25]. Our simulation also disgard the multiple scattering, which is recommended to be minimized in the real experimental setting by reducing the path length of the sample or by reducing the scattering contrast. Moreover, this framework can be applied to even more intricate systems, such as charged colloids, binary colloids[50], rods and charged rods[51], as well as other dense dispersion systems, where the strong coupling between the form factor and structure factor renders traditional analytic modeling ineffective. Since polydispersity is ubiquitous in nature, extending our framework to these cases will greatly improve the analytical power of the scattering technique.
@@ Acknowledgements.
We thank Chi-Huan Tung, Wei-Ren Chen, and Jan-Michael Carrillo for fruitful discussion. This research used resources at the Center for Nanophase Materials Sciences, US Department of Energy (DOE) Office of Science User Facilities operated by Oak Ridge National Laboratory. This research was sponsored by the Laboratory Directed Research and Development Program of Oak Ridge National Laboratory, managed by UT-Battelle, LLC, for the US DOE. Computations used resources of the Oak Ridge Leadership Computing Facility, which is supported by the DOE Office of Science under contract No. DE-AC05-00OR22725. Application of machine learning to soft matter was supported by the US DOE, Office of Science, Office of Basic Energy Sciences Data, Artificial Intelligence and Machine Learning at DOE Scientific User Facilities Program under award No. 34532.